# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

### v1

@@ Abstract
Spatial transcriptomics (ST) enables interrogating the molecular composition of tissue with ever-increasing resolution, depth, and sensitivity. However, costs, rapidly evolving technology, and lack of standards have constrained computational methods in ST to narrow tasks and small cohorts. In addition, the underlying tissue morphology as reflected by H&E-stained whole slide images (WSIs) encodes rich information often overlooked in ST studies. Here, we introduce HEST-1k, a collection of 1,108 spatial transcriptomic profiles, each linked to a WSI and metadata. HEST-1k was assembled using HEST-Library from 131 public and internal cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 cancer types. HEST-1k processing enabled the identification of 1.5 million expression–morphology pairs and 60 million nuclei. HEST-1k is tested on three use cases: (1) benchmarking foundation models for histopathology (HEST-Benchmark), (2) biomarker identification, and (3) multimodal representation learning. HEST-1k, HEST-Library, and HEST-Benchmark can be freely accessed via https://github.com/mahmoodlab/hest.
@@ 1 Introduction
Advances in molecular profiling enable spatially-resolved gene expression analysis with increasingly large gene panels, enhanced spatial resolution, and greater sensitivity Asp et al. (2020); Rao et al. (2021). From the early days of bulk RNA sequencing constrained by its coarse resolution and limited gene panels, spatially-resolved technologies have progressed to achieve whole-transcriptome sequencing at sub-cellular resolution Li and Wang (2021). In cancer research, spatial transcriptomics (ST) holds particular promise for characterizing the tumor microenvironment, a key element in understanding disease progression and treatment response Ståhl et al. (2016b); Marx (2021); de Visser and Joyce (2023); Xiao and Yu (2021). With the large amount of transcriptomics data generated by a single ST sample (e.g., $>$10 million transcripts are detected in a typical 10x Genomics Xenium assay), computational methods are often used to identify promising targets and biomarkers, such as based on gene expression clustering or gene-set enrichment analysis Rao et al. (2021).
@@
However, high costs and rapidly evolving technology have constrained computational methods to narrow tasks and data cohorts of only few patients Meylan et al. (2022); Janesick et al. (2023); Valdeolivas et al. (2023). Consequently, we observe a lack of standardized resources and unified formats for storing ST, which limits the development of deep learning models on a large scale Moses and Pachter (2022). In addition, the underlying tissue morphology, traditionally visualized in hematoxylin and eosin (H&E)-stained tissue sections (whole-slide images, WSIs), is often overlooked in ST studies, despite encoding valuable information. In particular, pairs of ST and WSI enable analyzing expression changes in their morphological context, which may facilitate the discovery of morphological biomarkers (e.g., changes in nuclear shape) that correspond to gene regulation patterns. In translational medicine, ST could be employed for identifying biomarkers, which are then morphologically validated and characterized by widely deployed clinical modalities such as H&E. Alternatively, pairs of ST and WSI can enable multimodal tissue representation learning for joint modeling of the morphomolecular signature of tissue at a scale and resolution beyond bulk RNA sequencing Chen et al. (2022). Finally, the development of “foundation models” for encoding histopathology image patches Wang et al. (2021); Filiot et al. (2023); Chen et al. (2024); Huang et al. (2023) has increased the need for new, diverse, and challenging benchmarks beyond diagnostic tasks. Using ST, new tasks can be defined for predicting gene expression from histology.
@@
Here, we introduce HEST-1k, a collection of paired ST and H&E-stained WSIs curated from public and internal cohorts (Figure 1.a). HEST-1k comprises 1,108 samples from 131 cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 subtypes. Processing all samples in HEST-1k resulted in 1.5 million expression–morphology pairs and 60 million detected nuclei. With new cohorts frequently made public, we also introduce the HEST-Library, a Python package for querying HEST-1k and assembling new legacy data (Figure 1.b). We highlight the potential of HEST-1k through three use cases: (i) benchmarking foundation models for histology using the HEST-Benchmark, a set of ten tasks (nine human cancer types and ten organs) for gene expression prediction from histology and evaluated on ten state-of-the-art models (Figure 1.c), (ii) a proof-of-concept demonstrating the use of HEST-1k for biomarker characterization (Figure 1.d), and (iii) a proof-of-concept for expression-guided fine-tuning of foundation models for histology (Figure 1.e).
@@
Figure 1: Overview of HEST. a. Description of HEST-1k, a dataset of $n$=1,108 paired Spatial Transcriptomics, H&E-stained whole-slide images and metadata. b. Overview of HEST-Library functionalities. c., d., e. Potential applications of HEST-1k include benchmarking foundation models for histology (c.), biomarker discovery (d.) and multimodal representation learning (e.).
@@ 2 Related work
Libraries for ST analysis. Libraries to process, visualize, and analyze ST have been built around two core pipelines: Scanpy Wolf et al. (2018) (and the Anndata format) in Python and Seurat Butler et al. (2018) in R. Scanpy has served as the foundation for several subsequent developments such as Squidpy Palla et al. (2022) for spatial data exploration at cellular-, gene-, and morphological-level, SpatialData Marconato et al. (2024) for multi-technology integration and deep learning interfacing, and STlearn Pham et al. (2023) for cell-cell interactions and spatiotemporal trajectory analyses. In R, Seurat Butler et al. (2018) has been consolidated with packages such as BayeSpace Zhao et al. (2021) for clustering and spot super-resolution, and Giotto suites Chen et al. (2023) for preprocessing, data integration, and visualization of multiple ST technologies. 10x genomics also includes software analytics through the Xenium and Visium Explorer pipelines for multimodal visualization, nuclear segmentation, and cell deconvolution. However, none of these pipelines were designed to handle the diversity of legacy data, where datasets can suffer from missing or incorrect data, such as alignment mismatches, incorrect pixel resolution, inconsistent image file formats, etc.
@@
Molecular profile prediction from H&E. Molecular profiling from histology images has been explored both at (1) slide-level to predict bulk molecular status/changes from a WSI, and at (2) patch-level to predict local molecular status/changes from regions-of-interest. (1) Slide-level profiling has been proposed to predict the presence of gene mutations Saldanha et al. (2023); Wagner et al. (2023); Kather et al. (2020); Echle et al. (2021); Fu et al. (2020); Wang et al. (2019); Loeffler et al. (2022), microsatellite instability Wagner et al. (2023); Kather et al. (2019b), and gene expression changes Schmauch et al. (2020); El Nahhas et al. (2024); He et al. (2020); Alsaafin et al. (2023), among others. The motivation is two-fold: First, patient screening to substitute or emulate costly clinical molecular assays, and second, to identify morphological correlates of molecular alterations for discovering novel biomarkers. Such studies can be conducted on large patient cohorts as they mainly rely on data generated by the routine clinical workflow (e.g., using TCGA cohorts with $>$11,000 cases from 33 cancer types). (2) With ST, several works have explored predicting expression changes from regions-of-interest He et al. (2020); Xie et al. (2023); Pang et al. (2021); Zhao et al. (2024); Rahaman et al. (2023); Monjo et al. (2022); Mondol et al. (2023); Chung et al. (2024); Wang et al. (2024). Due to limited cohort sizes (typically one to ten patients), transfer learning has become the norm using pretrained networks based on ConvNets He et al. (2020) or Vision Transformers Xie et al. (2023); Pang et al. (2021). Transcriptomic measurements are inherently noisy, hence the development of context-based methods that integrate global and local information from surrounding ST spots Chung et al. (2024); Zhao et al. (2024); Wang et al. (2024). While recent technologies offer near-single-cell resolution (such as Visium HD and Xenium), legacy assays operate at a more coarse resolution, which can be upsampled using super-resolution techniques Zhao et al. (2021); Bergenstråhle et al. (2022); Zhang et al. (2024). The potential clinical and research implications of such methods are still being explored, with HEST-1k potentially catalyzing their large-scale development.
@@
Foundation models in pathology. A fundamental task in computational pathology is to extract general-purpose embeddings of image patches (typically 256$\times$256-pixel regions) that can then be used for downstream tasks, such as diagnosis or prognosis prediction. To achieve this, self-supervised learning (SSL) has been extensively applied  Koohbanani et al. (2021); Filiot et al. (2023); Vorontsov et al. (2023); Wang et al. (2021); Kang et al. (2023); Ciga et al. (2022); Wang et al. (2022); Azizi et al. (2023); Chen et al. (2024); Lu et al. (2024), such as based on the DINOv2 framework Oquab et al. (2023). Models for general-purpose patch encoding are trained on increasingly large and more diverse patient cohorts with models growing in model size (e.g., Virchow uses a ViT-Huge trained on more than 1.5M WSIs Vorontsov et al. (2023)). As the number of such models rapidly increases, new, diverse, and challenging benchmarks are needed to replace or complement well-established tasks where performance has saturated. HEST-Benchmark aims to address this by offering a set of ten patch-level tasks for gene expression prediction from histology.
@@
Patch-level benchmarks in histopathology. Early task and dataset contributions in computational pathology revolved around classifying small regions of interest. Over the years, a variety of benchmarks have been established: In prostate cancer, Gleason grading at pixel- and patch-level has been widely explored, with public resources such as PANDA Bulten et al. (2022), AGCC Huo et al. (2022), DiagSet Koziarski et al. (2024), and SICAPv2 Silva-Rodríguez et al. (2020). In colorectal cancer, datasets have been proposed for tissue classification, such as HunCRC Pataki et al. (2022), UniToPatho Barbano et al. (2021), MHIST Wei et al. (2021), and CRC-100k Kather et al. (2019a). In breast cancer, morphological subtyping has been studied (e.g., for atypical ductal hyperplasia detection), such as BACH Aresta et al. (2019), BRACS Brancati et al. (2021), and BreakHis Spanhol et al. (2015), and for lymph node metastasis detection with Patch CAMELYON (pCAM) Veeling et al. (2018), respectively. However, the performance on many of these datasets has saturated; for instance, Gleason scoring reaches similar or better performance than pathologists Bulten et al. (2022), which limits objective comparisons of new methods and hinders well-informed model selection for developing better features. Instead, HEST-Benchmark provides a collection of diverse and challenging tasks that enable assessing the predictive capabilities of foundation models for histology.
@@ 3 HEST-1k Dataset
We present HEST-1k, a dataset of paired ST, H&E-stained WSIs, and metadata (Figure 1.a). To this end, we extracted all publicly available cohorts that provide ST with H&E-stained whole-slide images. Specifically, we harvested data from 10x Genomics public datasets (TENX) [^1] https://www.10xgenomics.com/datasets, Mendeley (MEND) [^2] https://data.mendeley.com/, Spatial-Research (SPA) [^3] https://www.spatialresearch.org/, Zenodo (ZEN) [^4] https://zenodo.org/, the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/gds/, GitHub [^6] https://github.com/, the Human Cell Atlas [^7] https://data.humancellatlas.org/, and internal data cohorts. A summary of all sources is provided in Appendix Table A1 with specifics in Appendix Table A2, A4, A6,A7,A8,A9, and A10.
@@ 3.1 Metadata
As Spatial Transcriptomics were not intended for large-scale computational research, they are provided in various formats (e.g., images can be in JPG or TIFF format, with or without cross-modal alignment files) and resolutions. We unified all data with comprehensive metadata with generic-, histology-, and expression-related descriptors for all samples. Generic: We provide the reference to the original publication, download link, year of publication, license, and sample species. Each sample is then categorized as either healthy, cancer, treated (which refers to a post-compound administration), genetically modified (mostly knock-out mouse samples), or other diseases (with an additional description). All cancer samples were unified using the OncoTree code, a taxonomy of cancer types provided by the Memorial Sloan Kettering Cancer Center [^8] https://oncotree.mskcc.org. Finally, we include the organ using the highest level of the OncoTree taxonomy as a reference. Expression: We include the number of genes and spots per sample, the spot resolution and spacing, the total number of reads, and the mean number of reads per spot. We additionally provide the transcriptomic technology (ST, Visum, Visium HD, or Xenium). Histology: We provide the image resolution (in $\mu$m/pixel) and magnification as 10$\times$ (1.15 to 0.8 $\mu$m/px), 20$\times$ (0.8 to 0.4 $\mu$m/px) and 40$\times$ (0.4 to 0.1 $\mu$m/px). All images with a resolution lower than 1.15 $\mu$m/px were discarded to ensure an acceptable image quality. In addition, we provide the image size at the highest resolution, and tissue preparation protocol (frozen or formalin-fixed paraffin-embedded, FFPE).

### v2

@@ Abstract
Spatial transcriptomics enables interrogating the molecular composition of tissue with ever-increasing resolution and sensitivity. However, costs, rapidly evolving technology, and lack of standards have constrained computational methods in ST to narrow tasks and small cohorts. In addition, the underlying tissue morphology, as reflected by H&E-stained whole slide images (WSIs), encodes rich information often overlooked in ST studies. Here, we introduce HEST-1k, a collection of 1,229 spatial transcriptomic profiles, each linked to a WSI and extensive metadata. HEST-1k was assembled from 153 public and internal cohorts encompassing 26 organs, two species (Homo Sapiens and Mus Musculus), and 367 cancer samples from 25 cancer types. HEST-1k processing enabled the identification of 2.1 million expression–morphology pairs and over 76 million nuclei. To support its development, we additionally introduce the HEST-Library, a Python package designed to perform a range of actions with HEST samples. We test HEST-1k and Library on three use cases: (1) benchmarking foundation models for pathology (HEST-Benchmark), (2) biomarker exploration, and (3) multimodal representation learning. HEST-1k, HEST-Library, and HEST-Benchmark can be freely accessed at https://github.com/mahmoodlab/hest.
@@ 1 Introduction
Advances in molecular profiling enable spatially-resolved gene expression analysis with increasingly large gene panels, enhanced spatial resolution, and greater sensitivity [asp2020spatially, rao2021exploring]. From the early days of bulk RNA sequencing constrained by its coarse resolution and limited gene panels, spatially-resolved technologies have progressed to achieve whole-transcriptome sequencing at sub-cellular resolution [li2021from]. In cancer research, spatial transcriptomics (ST) holds particular promise for characterizing the tumor microenvironment, a key element in understanding disease progression and treatment response [staahl2016visualization, marx2021method, devisser2023evolving, xiao2021tumor]. With the large amount of transcriptomics data generated by a single ST sample (e.g., $>$10 million transcripts are detected in a typical 10x Genomics Xenium assay), computational methods are often used to uncover promising biomarkers, such as employing clustering methods for cell phenotyping [rao2021exploring].
@@
However, high costs and rapidly evolving technology have constrained computational methods to narrow tasks and data cohorts of only a few patients [meylan2022tertiary, janesick2023high, valdeolivas2023charting]. Consequently, we observe a lack of standardized resources and unified formats for handling ST, which limits the development of deep learning models on a large scale [moses2022museum]. In addition, the underlying tissue morphology, traditionally visualized in hematoxylin and eosin (H&E)-stained tissue sections (whole-slide images, WSIs), is often overlooked in ST studies, despite encoding valuable information. In particular, pairs of ST and WSI enable analyzing expression changes in their morphological context, which may facilitate the identification of morphological biomarkers (e.g., changes in nuclear shape) that correspond to gene regulation patterns. Alternatively, pairs of ST and WSI can enable multimodal tissue representation learning for joint modeling of the morphomolecular signature of tissue at a scale and resolution beyond bulk RNA sequencing [chen2022pan]. Finally, the development of “foundation models” for encoding histopathology images [wang2021transpath, filiot2023scaling, chen2024towards, huang2023visual, lu2024towards] has increased the need for new, diverse, and challenging benchmarks beyond diagnostic tasks. Using ST, new tasks can be defined to predict gene expression changes from histology.
@@
Here, we introduce HEST-1k, a collection of paired ST and H&E-stained WSIs curated from public and internal cohorts (Figure 1.a). HEST-1k comprises 1,229 samples from 153 cohorts encompassing 26 organs, two species (Homo Sapiens and Mus Musculus), and 367 cancer samples from 25 different subtypes. Processing all samples in HEST-1k resulted in 2.1 million expression–morphology pairs and 76 million detected nuclei. With new cohorts frequently made public, we also introduce the HEST-Library, a Python package for interacting with HEST-1k data and assembling new samples as they become available (Figure 1.b). We highlight the potential of HEST-1k through three use cases: (i) benchmarking foundation models for histology using the HEST-Benchmark, a set of nine tasks (eight human cancer types and nine organs) for gene expression prediction from histology and evaluated on eleven state-of-the-art models (Figure 1.c), (ii) a proof-of-concept demonstrating the use of HEST-1k for biomarker characterization (Figure 1.d), and (iii) a proof-of-concept for expression-guided fine-tuning of foundation models for histology (Figure 1.e).
@@
Figure 1: The HEST environment. a. Overview of HEST-1k, a dataset of $n$=1,229 paired spatial transcriptomics, H&E-stained whole-slide images and metadata. “Pathological” cases refer to non-tumor/non-cancer samples; “Tumor” refers to non-cancer samples. b. Overview of HEST-Library functionalities. c., d., e. Applications of HEST-1k include benchmarking foundation models for histology (c.), biomarker exploration (d.) and multimodal representation learning (e.).
@@ 2 Related work
Libraries for ST analysis. Libraries to process, visualize, and analyze ST have been built around two core pipelines: Scanpy [wolf2018scanpy] (and the Anndata format) in Python and Seurat [butler2018integrating] in R. Scanpy has served as the foundation for several subsequent developments such as Squidpy [palla2022squidpy] for spatial data exploration at cellular-, gene-, and morphological-level, SpatialData [marconato2024spatialdata] for multi-technology integration and deep learning interfacing, STlearn [pham2023robust] for cell-cell interactions and spatiotemporal trajectory analyses, and SOPA [blampey2024sopa] for designing multistep pipelines. In R, Seurat [butler2018integrating] has been consolidated with packages such as BayeSpace [zhao2021spatial] for clustering and spot super-resolution, and Giotto suites [chen2023giotto] for preprocessing, data integration and visualization of multiple ST technologies. 10x Genomics also includes proprietary software analytics through the Xenium and Visium Explorer pipelines for multimodal visualization, nuclear segmentation, and cell deconvolution. However, none of these pipelines were designed to handle the diversity of legacy data, where datasets can suffer from missing or incorrect data, such as alignment mismatches, incorrect pixel resolution, inconsistent image file formats, etc.
@@
Molecular profile prediction from H&E. Molecular profiling from histology images has been explored both at (1) slide-level to predict bulk molecular status/changes from a WSI and at (2) patch-level to predict local molecular status/changes from regions-of-interest. (1) Slide-level profiling has been explored to predict the gene mutations [saldanha2023self, wagner2023fully, kather2020pan, echle2021deep, fu2020pan, wang2019predicting, loeffler2022artificial], microsatellite instability [wagner2023fully, kather2019deep], and gene expression changes [schmauch2020deep, el2024regression, he2020integrating, alsaafin2023learning], among others. The motivation is two-fold: First, patient screening to substitute or complement costly clinical molecular assays, and second, to identify morphological correlates of molecular alterations for discovering novel biomarkers. Such studies can be conducted on large patient cohorts as they mainly rely on data generated by the routine clinical workflow (e.g., using TCGA cohorts with $>$11,000 cases from 33 cancer types). (2) With ST, several works have explored predicting expression changes from regions-of-interest [he2020integrating, xie2023spatially, pang2021leveraging, zhao2024hist2cell, rahaman2023breast, monjo2022efficient, mondol2023hist2rna, chung2024accurate, wang2024m2ort]. Due to limited cohort sizes (typically one to ten patients), transfer learning has become the norm using pretrained models based on ConvNets [he2020integrating] or Vision Transformers [xie2023spatially, pang2021leveraging]. Due to the inherent noise found in transcriptomic measurements, several methods have been developed for integrating context that can account for global and local information from surrounding ST spots [hu2021spagcn, chung2024accurate, zhao2024hist2cell, wang2024m2ort]. While recent technologies offer near-single-cell resolution (such as Visium HD and Xenium), legacy assays operate at a more coarse resolution, which can be upsampled using super-resolution techniques [zhao2021spatial, bergenstrahle2022super, zhang2024inferring]. The potential clinical and research implications of such methods are still being explored, with HEST-1k potentially catalyzing their large-scale development.
@@
Foundation models in pathology. A fundamental task in computational pathology is to extract general-purpose embeddings of image patches (typically 256$\times$256 to 512$\times$512-pixel regions) that can then be used for downstream tasks, such as diagnosis or prognosis prediction. To achieve this, self-supervised learning (SSL) has been extensively applied [koohbanani2021self, filiot2023scaling, vorontsov2024foundation, wang2021transpath, kang2023benchmarking, ciga2022self, wang2022transformer, azizi2023robust, chen2024towards, lu2024towards, jaume2024transcriptomics, jaume2024multistain], such as based on the DINOv2 framework [oquab2023dinov2]. General-purpose patch encoders are trained on increasingly large and diverse patient cohorts (e.g., UNI [chen2024towards] uses a ViT-Large trained on 100k WSIs, Virchow [vorontsov2024foundation] uses a ViT-Huge trained on more than 1.5M WSIs). Recently, vision-language encoders designed for pathology have also been proposed [gamper2021multiple, lu2023visual, huang2023visual, lu2024towards, lu2024multimodal] and rely on large-scale paired data scraped from social media, textbooks, or publications. As the number of such models rapidly increases, new, diverse, and challenging benchmarks are needed to replace or complement well-established tasks where performance has saturated. HEST-Benchmark aims to address this by offering a set of nine patch-level tasks for gene expression prediction from histology.
@@
Patch-level benchmarks in histopathology. Early task and dataset contributions in computational pathology revolved around classifying small regions of interest. Over the years, a variety of benchmarks have been established: In prostate cancer, Gleason grading at pixel- and patch-level has been widely explored, with public resources such as AGGC [huo2022comprehensive], DiagSet [koziarski2024diagset], and SICAPv2 [silvarodriguez2020going]. In colorectal cancer, datasets have been proposed for tissue classification, such as HunCRC [pataki2022huncrc], UniToPatho [barbano2021unitopatho], MHIST [wei2021petri], and CRC-100k [kather2019predicting]. In breast cancer, morphological subtyping has been vastly explored (e.g., for atypical ductal hyperplasia detection), such as BACH [aresta2019bach], BRACS [brancati2021bracs], and BreakHis [spanhol2015dataset], and for lymph node metastasis detection with Patch CAMELYON (pCAM) [veeling2018rotation], respectively. However, the performance on many of these datasets has saturated; for instance, Gleason scoring reaches similar or better performance than pathologists [bulten2022artificial], which limits objective comparisons of new methods and hinders well-informed model selection for developing better features. Instead, HEST-Benchmark provides a collection of diverse and challenging tasks that enable assessing the predictive capabilities of foundation models for histology.
@@ 3 HEST-1k Dataset
We present HEST-1k, a dataset of paired ST, H&E-stained WSIs, and metadata (Figure 1.a). To this end, we extracted all publicly available cohorts that provide ST with H&E-stained whole-slide images. Specifically, we harvested data from 10x Genomics public datasets (TENX) [^1] https://www.10xgenomics.com/datasets, Mendeley (MEND) [^2] https://data.mendeley.com/, Spatial-Research (SPA) [^3] https://www.spatialresearch.org/, Zenodo (ZEN) [^4] https://zenodo.org/, the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/gds/, GitHub [^6] https://github.com/, the Human Cell Atlas [^7] https://data.humancellatlas.org/, BioStudies [^8] https://www.ebi.ac.uk/biostudies/, HTAN [^9] https://humantumoratlas.org/, and internal data cohorts. A summary of all sources is provided in Appendix Table A1 with specifics in Appendix Table A2, A4, A6,A7,A8,A9, and A10.
@@ 3.1 Metadata
As spatial transcriptomics experiments were not intended for large-scale computational research, they are provided in various formats (e.g., images can be in JPG or TIFF format, with or without cross-modal alignment files) and resolutions. We unified all data with comprehensive metadata with generic-, histology-, and expression-related descriptors for all samples. Generic: We provide the reference to the original publication, download link, year of publication, license, and sample species. Each sample is then categorized as either healthy, cancer, tumor (non-cancer), treated (which refers to a post-compound administration), genetically modified (mostly knock-out mouse samples), or pathological (i.e., non-tumorous with extra specification). All cancer samples were unified using the OncoTree code, a taxonomy of cancer types provided by the Memorial Sloan Kettering Cancer Center [^10] https://oncotree.mskcc.org. Finally, we provide the organ using the highest level of the OncoTree taxonomy as a reference. Expression: We report the number of genes and spots per sample, the spot resolution and spacing, the total number of reads, and the mean number of reads per spot. We additionally provide the transcriptomic technology (ST, Visum, Visium HD, or Xenium). Histology: We provide the image resolution (in $\mu$m/pixel) and magnification as 10$\times$ (1.15 to 0.8 $\mu$m/px), 20$\times$ (0.8 to 0.4 $\mu$m/px) and 40$\times$ (0.4 to 0.1 $\mu$m/px). All images with a pixel size higher than 1.15 $\mu$m/px were discarded to ensure an acceptable image quality. In addition, we provide the image size at the highest resolution and the tissue preparation protocol (frozen or formalin-fixed paraffin-embedded, FFPE).